### v1

@@ A.2 Training
Full configurations are listed in Tables 6 and 7. Training ran on $4\times$ NVIDIA A100 (80 GB) or $8\times$ NVIDIA RTX A6000 (48 GB) GPUs; each epoch took roughly $30$ minutes for 1515k trajectories. We used per-device batch sizes of 1122 with gradient accumulation to keep the effective batch size fixed at $64$.
@@ A.3 Evaluation
We use lm_eval (EleutherAI’s lm-eval-harness) for dataset loading, prompting, and scoring. We apply the harness’s default post-processors. Decoded text is truncated at any task-specific stop_sequences prior to scoring. For HumanEval and MBPP, we additionally employ Python post-processing scripts adapted from the dParallel repository chen2025dparallel to handle code execution and pass@1 computation.
@@
Appendix B Limitations and Future Directions
@@
Our current training relies on offline, static trajectories. Although we carefully select datasets to limit domain bias, the student can still overfit to the teacher’s priors because supervision does not adapt during training. A straightforward direction is to broaden the data mixture and scale beyond our current $\sim$15k samples, covering more domains.
@@
A complementary path is to move from offline supervision to on-the-fly generation. If the student generates trajectories during training and the teacher verifies or refines them online, we can incorporate on-policy learning to reduce the training-inference discrepancy. In practice, this is currently constrained by DLM sampling speed: slow generation limits the feasibility of closed-loop training without prior trajectory materialization. Improving DLM inference throughput is therefore a prerequisite for large-scale online training.
@@
CDLM’s performance is ultimately bounded by the teacher: a bidirectional DLM distilled into a block-causal student cannot exceed the teacher’s knowledge. To lift this ceiling, distillation from stronger autoregressive (AR) teachers is a natural next step. In this work we intentionally used high-quality, third-party-curated Qwen2.5-7B generations lansechen_easy_2025; lansechen_hard_2025 as ground-truth text, hoping to inject AR knowledge into the student. A more direct, and likely more effective, approach is to distill from an AR teacher: AR models are already scaled up, typically stronger, and much faster at generation than naive DLMs, making large-scale dataset construction easier. Moreover, distillation is well suited to transferring knowledge from a larger, more powerful model to smaller students (here, Dream-7B and LLaDA-8B).

### v2

@@ A.2 Training
Training Configurations.
Models are optimized with AdamW using a constant learning-rate schedule with warmup. Prompts are left-padded to $512$ tokens and decoded with generation length $256$, resulting in a fixed sequence length of $768$. Based on the ablation study in Section 5.3.1, we set the loss weights $(w_{\text{distill}},w_{\text{cons}},w_{\text{dlm}})$ to (1.0, 0.5, 0.01)(1.0,\,0.5,\,0.01) for Dream and (1.0, 0.5, 0.1)(1.0,\,0.5,\,0.1) for LLaDA. All models are trained for $16$ epochs with an effective batch size of $64$, and we select the best checkpoint based on validation performance. Training ran on $4\times$ NVIDIA A100 (80 GB) or $8\times$ NVIDIA RTX A6000 (48 GB) GPUs; each epoch took roughly $30$ minutes for 1515k trajectories. We used per-device batch sizes of 1122 with gradient accumulation to keep the effective batch size fixed at $64$. Full configurations are listed in Tables 5 and 6.
@@ A.3 Evaluation
We use the lm_eval framework (EleutherAI’s lm-eval-harness) for dataset loading, prompting, and evaluation. We apply the harness’s default post-processing rules, truncating decoded outputs at any task-specific stop_sequences prior to scoring. Evaluation metrics include exact match scores on GSM8K and math-verify scores on MATH. For code-generation benchmarks (HumanEval and MBPP), we additionally employ Python post-processing scripts adapted from the DParallel repository Chen et al. (2025) to execute generated code and compute pass@1.
@@
Inference settings.
For LLaDA, we enable block-wise decoding with block size B=32B{=}32. Dream does not natively support block-wise decoding, so we extend the baseline implementation to support block-wise decoding under the same configuration.
Appendix B Additional Results
Figure 7: Validation trends during CDLM–Dream training. Validation score (red) and average refinement iterations (blue) across training epochs on GSM8K-CoT (8-shot, left) and MBPP-Instruct (0-shot, right).
Figure 8: Impact of inference-time block size. Accuracy and throughput (TPS) across different block sizes $B\in\{4,8,16,32,64\}$ for CDLM–Dream and CDLM–LLaDA. All models are trained with B=32B{=}32 and evaluated by varying the inference-time block size.
B.1 Validation Trends During Training
We track validation performance and average refinement iterations across training epochs for CDLM–Dream, following the training setup in Table 5. Figure 7 shows the trends on GSM8K-CoT (left) and MBPP-Instruct (right). On GSM8K-CoT, validation accuracy steadily improves in early epochs and then saturates, while the average number of refinement iterations consistently decreases. This indicates that training encourages the model to reach correct solutions in fewer steps, aligning with our goal of accelerating inference. On MBPP-Instruct, validation accuracy exhibits moderate fluctuations and a downward trend in later epochs, whereas the average number of refinement iterations continues to decrease. This suggests that continued training further reduces the number of refinement iterations, but may also bias the model toward math-style reasoning due to the composition of the training data. We therefore select checkpoints before overfitting becomes pronounced.
Table 7: Ablation of the confidence threshold. Performance under different token-level confidence thresholds $\tau_{\text{conf}}$ for CDLM–Dream on GSM8K-CoT (8-shot) and HumanEval-Instruct (0-shot).
GSM8K -CoT (8-shot) 0.95 42.7 2.5 78.8
0.90 51.7 2.1 78.8
0.85 57.7 1.9 78.4
HumanEval -Instruct (0-shot) 0.95 34.4 2.8 51.2
0.90 43.3 2.2 50.0
0.85 47.3 2.0 48.2
B.2 Token-Level Confidence Threshold
B.3 Sensitivity Analysis on Inference Block Size
We study the effect of varying the inference-time block size by evaluating CDLM models trained with B=32B{=}32 under block sizes $B\in\{4,8,16,32,64\}$. Figure 8 reports accuracy and throughput (TPS) on representative benchmarks.
Across both CDLM–Dream and CDLM–LLaDA, throughput increases monotonically as the block size grows from B=4B{=}4 to B=32B{=}32, reflecting more aggressive exploitation of parallelism and a gradual move away from AR-style decoding. However, increasing the block size beyond the training configuration, from B=32B{=}32 to B=64B{=}64, does not consistently yield further speedups. Instead, it often results in saturated or even reduced TPS, indicating a train–inference discrepancy because the models are not trained to predict and complete such large blocks.
For most tasks, the best accuracy is achieved when the inference-time block size matches the training block size (B=32B{=}32), suggesting that aligning training and inference configurations helps avoid performance degradation. To support a wider range of block sizes at inference time, future work could explore curriculum-based training schemes that begin with smaller block sizes (e.g., B=8B{=}8 or $16$) and gradually increase the difficulty of completing multiple blocks.
Figure 9: Roofline analysis of batched inference across language model variants. Roofline placement of decoding in autoregressive (AR) models, vanilla DLMs, and block-wise DLMs with block sizes $B\in\{4,16,32\}$ under batch sizes $\mathrm{bs}\in\{1,2,4,8,16,32,64,128\}$. All results are obtained via roofline simulation assuming an NVIDIA A100-SXM4-80GB GPU with peak FP16 performance of 311.9 TFLOP/s and memory bandwidth of 2039.0 GB/s.
B.4 Roofline Analysis of AR Models, Vanilla DLMs, and Block-wise DLMs
Following Section 5.4, we also perform a roofline analysis to characterize the hardware efficiency of AR decoding, full-attention vanilla DLMs, and block-wise DLMs (i.e., CDLM). We derive roofline parameters for an NVIDIA A100-SXM4-80GB (GA100) GPU NVIDIA (2020; 2022), assuming dense FP16 Tensor Core computation (without 2:4 structured sparsity) at the boost clock frequency. The theoretical peak FP16 performance is
[…]
With memory bandwidth of $2039.0~\mathrm{GB/s}$, the theoretical ridge point is
$\mathrm{AI}_{\mathrm{ridge}}=\frac{311.9~\mathrm{TFLOP/s}}{2039.0~\mathrm{GB/s}}\approx 153.0~\mathrm{FLOP/Byte}.$
Note that the AR and DLM backbones differ (e.g., GQA vs. MHA), so absolute AI values are not directly comparable across model families.
AR Models and Vanilla DLMs.
In Figure 9 (left), AR models start in a strongly memory-bound regime, with AI near 1 at $\mathrm{bs}=1$. Even at $\mathrm{bs}=128$, AI remains well below the ridge point, and performance scales noticeably with batch size. In contrast, vanilla DLMs are compute-bound even at $\mathrm{bs}=1$, with AI ($\approx 438.9$) far exceeding the ridge point. Accordingly, performance quickly saturates and shows little improvement with increasing batch size. The observed plateau lies slightly below the theoretical peak, as non-linear operations such as layer normalization and softmax are executed on vector units with lower peak throughput than Tensor Cores. This heterogeneity in execution units results in a lower effective performance ceiling during inference.
Block-wise DLMs.
Figure 9 (right) shows the block-wise DLMs with block sizes $B=4,16,32$. At $\mathrm{bs}=1$, their AI values (4.0, 15.8, and 31.1, respectively) are higher than AR decoding but lower than vanilla DLMs. This reflects increased intra-block data reuse: processing multiple tokens within a shared KV-cache context and reusing the same model weights across tokens in the block amortizes memory access and raises AI. For small batch sizes, block-wise DLMs remain memory-bound and exhibit strong scaling as batch size increases. As $B$ grows, performance tends to saturate around $\mathrm{bs}=64$ for $B=4$, $\mathrm{bs}=16$ for $B=16$, and $\mathrm{bs}=8$ for $B=32$.
Overall, block-wise DLMs occupy a middle ground between AR models, which are memory-bound, and vanilla DLMs, which are compute-bound. By exploiting parallelism within blocks while retaining moderate AI, block-wise decoding improves compute utilization in low-batch regimes without immediately saturating available compute.
Appendix C Limitations and Future Directions
@@
Our current training relies on offline, static trajectories. Although we carefully select datasets to limit domain bias, the student can still overfit to the teacher’s priors because supervision does not adapt during training. A straightforward direction is to broaden the data mixture and scale beyond our current $\sim$15k samples to cover more diverse domains.
@@
A complementary path is to move from offline supervision to on-the-fly generation. If the student generates trajectories during training and the teacher verifies or refines them online, on-policy learning could reduce the training–inference discrepancy. In practice, this is currently constrained by DLM sampling speed: slow generation limits the feasibility of closed-loop training without prior trajectory materialization. Improving DLM inference throughput is therefore a prerequisite for large-scale online training.
Another promising direction is to use CDLM as a draft model within speculative decoding frameworks Leviathan et al. (2023), where diffusion models can serve as fast proposers for autoregressive verifiers. A naive DLM drafter yields limited benefit because standard DLMs require many refinement steps to produce coherent text. In contrast, a consistency-trained DLM (CDLM) can generate reasonable drafts in far fewer steps. Pairing a CDLM drafter with an autoregressive (AR) verifier (cf. Christopher et al. (2025)) is promising: the CDLM proposes multi-token candidates quickly, and the AR model supplies strong final validation.
@@
CDLM’s performance is ultimately bounded by the teacher: a bidirectional DLM distilled into a block-causal student cannot exceed the teacher’s knowledge. To lift this ceiling, distillation from stronger autoregressive (AR) teachers is a natural next step. In this work we intentionally used high-quality, third-party-curated Qwen2.5-7B generations Lansechen (2025a; b) as ground-truth text, hoping to inject AR knowledge into the student. A more direct, and likely more effective, approach is to distill from an AR teacher: AR models are already scaled up, typically stronger, and much faster at generation than naive DLMs, making large-scale dataset construction easier. Identifying effective ways to extract suitable training trajectories from AR models is therefore a key prerequisite.